OpenStudy (anonymous):

Algebra 2?!?!?!?! WILL GIVE MEDAL FOR BEST ANSWER AND FAN!

OpenStudy (anonymous):

Which of the following functions represents the following graph? f(x) = the square root of x and 2 – 1 f(x) = (x + 2)^(2) – 1 f(x) = negative 1 over x plus 2 f(x) = (x + 2)^(3) – 1

OpenStudy (anonymous):

Notice the asymptote at -2. This should help determine which equation is the answer. Keep in mind that an asymptote is anywhere a function is not defined. A function cannot be defined when the denominator is 0.

OpenStudy (anonymous):

There are also other times where functions aren't defined or certain values of x are not in their domain. Like square roots cannot have a negative radican, logarithmic functions cannot have a negative parameter, etc.

OpenStudy (anonymous):

Ok..

OpenStudy (anonymous):

Do D is not the right answer, right?

OpenStudy (anonymous):

Right. You could put any number into x and it would still be valid. I said a function cannot be defined when the DENOMINATOR is 0. There's only one with a denominator.

OpenStudy (anonymous):

C. has a denominator

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

I tagged you in it